OpenStudy (anonymous):

Indicate in standard form the equation of the line passing through the given points. L(5, 0), M(0, 5)

OpenStudy (anonymous):

@Zweifel

OpenStudy (anonymous):

@dan815

OpenStudy (acxbox22):

first find the slope

OpenStudy (anonymous):

-5/5

OpenStudy (acxbox22):

which is...?

OpenStudy (anonymous):

equal to -1

OpenStudy (acxbox22):

and we have our y intercept so we can put it in slope intercept form y=-x+5 now convert to standard form by isolating the x and y on the left

OpenStudy (anonymous):

y-x=5?

OpenStudy (anonymous):

@acxbox22

OpenStudy (anonymous):

@dan815

OpenStudy (acxbox22):

you add x to both sides so it is x+y=5

OpenStudy (anonymous):

oh sorry didn't see the negative ok we've got that now what

OpenStudy (anonymous):

nevermind I figured it out that right there is actually my answer Thank you

OpenStudy (acxbox22):

np :)